OpenStudy (anonymous):

which set of fractions is ordered from least to greatest? 7/8,5/11,2/3

OpenStudy (cwrw238):

convert each fraction so its denominator LCM of 8 , 11 and 3 which is 8*11*3 = 264

OpenStudy (cwrw238):

so 7 / 8 = 7 * 11*3 = 231 ------ --- 264 264

OpenStudy (cwrw238):

convert the other 2 and you'll easily see the order you require